OpenStudy (blackops2luvr):

The formula r = d/t gives the rate r for a distance d traveled in an amount of time t. A trip of 120 miles takes 2.5 hours.    What was the rate for this trip?   A. 48 mi/hr   B. 60.5 mi/hr   C. 122.5 mi/hr   D. 300 mi/hr

OpenStudy (blackops2luvr):

any ideas?

OpenStudy (cggurumanjunath):

r = d/t=120/2.5